# Global Existence for General Systems of Isentropic Gas Dynamics via a Weighted Pressure Perturbation Approach

### v1

@@ 4.1 Evolution of Riemann Invariants
4.2 Invariant Regions and $L^{\infty}$ Bounds
We now establish the uniform $L^{\infty}$ bounds for the approximate viscous solutions. Define the phase–plane invariant region:
$\Sigma=\big\{(\rho,u):\rho\geq\delta,\quad z(\rho,u)\geq C_{\inf},\quad w(\rho,u)\leq C_{\sup}\big\},$ (4.6)
where $C_{\inf}=\inf_{x}z(\rho_{0}(x),u_{0}(x))$ and $C_{\sup}=\sup_{x}w(\rho_{0}(x),u_{0}(x))$.
Theorem 4.1 (Invariant Regions).
Assume $(\rho_{0},u_{0})\in\Sigma$. Then for any $\varepsilon>0$, the solution $(\rho^{\varepsilon},u^{\varepsilon})$ of the viscous system (4.1) satisfies
$(\rho^{\varepsilon}(x,t),u^{\varepsilon}(x,t))\in\Sigma\quad\text{for all }t>0.$
@@ 4.2 Invariant Regions and $L^{\infty}$ Bounds
The key is the sign of $K(\rho)$ in (4.3). Since $K(\rho)>0$ for all $\rho\geq\delta$, the quadratic term satisfies $\varepsilon K(\rho)\rho_{x}^{2}\leq 0$.
Step 1: Maximum principle for $w$. At any interior point $(x_{0},t_{0})$ where $w$ attains a local maximum, we have $w_{t}\geq 0$, $w_{x}=0$, $w_{xx}\leq 0$. Equation (4.3) implies $w_{t}\leq 0$. By the standard parabolic maximum principle, $w(x,t)\leq C_{\sup}$.
Step 2: Minimum principle for $z$. Consider the evolution equation for $z$ derived in (4.4):
$z_{t}+\lambda_{1}z_{x}=\varepsilon z_{xx}+\frac{2\varepsilon}{\rho}\rho_{x}z_{x}+\varepsilon K(\rho)\rho_{x}^{2}.$
Let $(x_{0},t_{0})$ be a point where $z$ attains a local minimum in space. At this point, $z_{x}(x_{0},t_{0})=0$ and $z_{xx}(x_{0},t_{0})\geq 0$. Evaluating the equation at this minimum yields:
$z_{t}=\varepsilon z_{xx}+\varepsilon K(\rho)\rho_{x}^{2}.$
Since $\varepsilon>0$, $z_{xx}\geq 0$, and $K(\rho)>0$ for $\rho\geq\delta$, the right-hand side is non-negative. Thus, $z_{t}(x_{0},t_{0})\geq 0$. This implies that the minimum value of $z$ is non-decreasing with respect to time. Consequently, $z$ cannot fall below its initial minimum:
$z(x,t)\geq\inf_{y}z(\rho_{0}(y),u_{0}(y))=C_{\inf},\quad\forall t>0.$
Step 3: Lower bound $\rho^{\varepsilon}\geq\delta$. We consider the difference $w-z=2\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds=:F(\rho)$. Since $F(\rho)$ is strictly increasing for $\rho\geq\delta$ with $F(\delta)=0$, the condition $\rho\geq\delta$ is equivalent to $w\geq z$. Initially, $w_{0}\geq z_{0}$ since $\rho_{0}\geq\delta$. The evolution equations for $w$ and $z$ both contain the negative source term $\varepsilon K(\rho)\rho_{x}^{2}$. Subtracting the equation for $z$ from $w$ preserves the non-negativity of wzw-z due to the favorable sign of the dissipation. Thus, $\rho^{\varepsilon}(x,t)\geq\delta$.
Step 4: Upper bound on $\rho^{\varepsilon}$. From the bounds on $z$ and $w$:
$\int_{\delta}^{\rho^{\varepsilon}(x,t)}\frac{\tilde{c}(s)}{s}\,ds=\frac{w-z}{2}\leq\frac{C_{\sup}-C_{\inf}}{2}.$
Assumption (1.5) implies that this integral diverges as $\rho\to\infty$. Therefore $\rho^{\varepsilon}$ is uniformly bounded.
Step 5: Bound for $u^{\varepsilon}$. Since $u=(w+z)/2$, the bounds on $w,z$ imply uniform boundedness for $u^{\varepsilon}$. ∎
@@ 4.3 Energy Dissipation and $H^{-1}$ Compactness
$D^{\varepsilon}:=\varepsilon(\mathbf{U}^{\varepsilon}_{x})^{T}\nabla^{2}\eta^{*}\mathbf{U}^{\varepsilon}_{x}=\varepsilon\Big(\frac{\tilde{P}^{\prime}(\rho^{\varepsilon})}{(\rho^{\varepsilon})^{2}}(\rho^{\varepsilon}_{x})^{2}+\rho^{\varepsilon}(u^{\varepsilon}_{x})^{2}\Big)\geq 0.$

### v2

@@ 4.1 Evolution of Riemann Invariants
4.2 Invariant Regions via the Chueh-Conley-Smoller Theory
Now we establish uniform $L^{\infty}$ bounds by appealing to the general theory of positively invariant regions for systems of nonlinear parabolic equations, as developed by Chueh, Conley, and Smoller [8]. This framework provides a rigorous geometric justification for the bounds in the presence of coupled diffusion and convection.
Consider the viscous system written in terms of the conservative variables $\mathbf{U}=(\rho,m)^{T}$ with $m=\rho u$:
$\mathbf{U}_{t}+\mathbf{F}(\mathbf{U})_{x}=\varepsilon\mathbf{U}_{xx},$ (4.6)
where the flux is $\mathbf{F}(\mathbf{U})=(m,m^{2}/\rho+\tilde{P}(\rho))^{T}$ and the diffusion matrix is the scalar matrix $D=\varepsilon I_{2}$.
We define the phase-space region $\Sigma\subset\mathbb{R}^{2}$ as the intersection of the regions defined by the Riemann invariants and the cut-off density:
$\Sigma=\left\{\mathbf{U}\in\mathbb{R}^{2}:z(\mathbf{U})\geq C_{\inf},\quad w(\mathbf{U})\leq C_{\sup},\quad\rho\geq\delta\right\},$ (4.7)
where $C_{\inf}=\inf_{x}z(\mathbf{U}_{0}(x))$ and $C_{\sup}=\sup_{x}w(\mathbf{U}_{0}(x))$. Note that the third condition $\rho\geq\delta$ is geometrically redundant if the integral $\int_{\delta}^{\rho}\frac{\tilde{c}}{s}ds$ spans the full range, but strictly speaking, it represents the degeneracy boundary $w(\mathbf{U})-z(\mathbf{U})=0$.
Lemma 4.1 (Convexity of Invariant Regions).
Suppose the perturbed pressure satisfies the genuine nonlinearity condition $2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)>0$ for $\rho>\delta$. Then, the phase-space region $\Sigma$ defined by the intersection of the sets:
$\mathcal{R}_{w}=\{(\rho,m):w(\rho,m)\leq C_{\sup}\},\quad\mathcal{R}_{z}=\{(\rho,m):z(\rho,m)\geq C_{\inf}\},\quad\mathcal{R}_{\rho}=\{(\rho,m):\rho\geq\delta\}$
is a convex subset of $\mathbb{R}^{2}$.
@@ 4.2 Invariant Regions via the Chueh-Conley-Smoller Theory
The region $\mathcal{R}_{\rho}$ is a half-plane $\rho\geq\delta$, which is trivially convex. To prove the convexity of $\mathcal{R}_{w}$ and $\mathcal{R}_{z}$, we examine the curvature of their boundary curves in the $(\rho,m)$-plane.
1. Convexity of $\mathcal{R}_{w}$: The boundary of $\mathcal{R}_{w}$ is the level curve $w(\rho,u)=C_{\sup}$. Expressing $m=\rho u$ as a function of $\rho$ along this curve:
$m_{w}(\rho)=\rho u=\rho\left(C_{\sup}-\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds\right).$
Differentiating with respect to $\rho$, and noting that along this curve $du=-\frac{\tilde{c}}{\rho}d\rho$:
$m_{w}^{\prime}(\rho)=u+\rho u^{\prime}=u+\rho\left(-\frac{\tilde{c}}{\rho}\right)=u-\tilde{c}$.
We compute the second derivative to determine the curvature:
$m_{w}^{\prime\prime}(\rho)=\frac{d}{d\rho}(u-\tilde{c})=u^{\prime}-\tilde{c}^{\prime}(\rho)=-\frac{\tilde{c}}{\rho}-\tilde{c}^{\prime}(\rho)=-\frac{1}{\rho}(\tilde{c}(\rho)+\rho\tilde{c}^{\prime}(\rho)).$
Recall that $\tilde{c}^{2}(\rho)=\tilde{P}^{\prime}(\rho)$. Differentiating gives $2\tilde{c}\tilde{c}^{\prime}=\tilde{P}^{\prime\prime}$. Thus $\tilde{c}^{\prime}=\frac{\tilde{P}^{\prime\prime}}{2\tilde{c}}$. Substituting this back:
$\tilde{c}+\rho\tilde{c}^{\prime}=\tilde{c}+\frac{\rho\tilde{P}^{\prime\prime}}{2\tilde{c}}=\frac{2\tilde{c}^{2}+\rho\tilde{P}^{\prime\prime}}{2\tilde{c}}=\frac{2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}}{2\tilde{c}}$.
Since $\tilde{c}>0$ for $\rho>\delta$ and the genuine nonlinearity condition $2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}>0$ holds (as proved in Lemma 2.1), we have $\tilde{c}+\rho\tilde{c}^{\prime}>0$. Consequently, $m_{w}^{\prime\prime}(\rho)<0$. This implies that the function $m_{w}(\rho)$ is strictly concave. The sublevel set $\mathcal{R}_{w}$ (the region ”below” this concave curve) is therefore a convex set.
2. Convexity of $\mathcal{R}_{z}$: Similarly, the boundary of $\mathcal{R}_{z}$ is $z(\rho,u)=C_{\inf}$, which implies $u=C_{\inf}+\int\frac{\tilde{c}}{s}$. The slope of the boundary curve $m_{z}(\rho)$ is:
$m_{z}^{\prime}(\rho)=u+\rho u^{\prime}=u+\rho\left(\frac{\tilde{c}}{\rho}\right)=u+\tilde{c}$.
The second derivative is:
$m_{z}^{\prime\prime}(\rho)=u^{\prime}+\tilde{c}^{\prime}=\frac{\tilde{c}}{\rho}+\tilde{c}^{\prime}=\frac{1}{\rho}(\tilde{c}+\rho\tilde{c}^{\prime})>0.$
Since $m_{z}^{\prime\prime}(\rho)>0$, the boundary curve is strictly convex. The superlevel set $\mathcal{R}_{z}$ (the region ”above” this convex curve) is therefore a convex set.
Since the intersection of convex sets is convex, $\Sigma=\mathcal{R}_{w}\cap\mathcal{R}_{z}\cap\mathcal{R}_{\rho}$ is a convex region in the conservative phase space. ∎
Theorem 4.2 (Positively Invariant Regions).
Assume the initial data $\mathbf{U}_{0}(x)$ lies in $\Sigma$ for all $x\in\mathbb{R}$. Then, the solution $\mathbf{U}(x,t)$ of the viscous system (4.6) remains in $\Sigma$ for all $t>0$. Consequently, the density satisfies the uniform lower bound $\rho(x,t)\geq\delta$.
Proof.
According to the foundational theory of Chueh, Conley, and Smoller [8] (Theorem 2.1 and Corollary 3.2), a region $\Sigma$ defined by a set of inequalities $\{G_{i}(\mathbf{U})\leq 0\}_{i=1}^{k}$ is positively invariant for the system (4.6) if the following conditions hold on the boundary $\partial\Sigma$:
  1. 1.
    Eigenvector Condition: At any boundary point $\mathbf{U}^{*}$ where $G_{i}(\mathbf{U}^{*})=0$, the gradient $\nabla_{\mathbf{U}}G_{i}(\mathbf{U}^{*})$ is a left eigenvector of the diffusion matrix $D$.
  2. 2.
    Quasi-Convexity: The region $\Sigma$ is convex (or the functions $G_{i}$ are quasi-convex).
  3. 3.
    Flux Condition (Characteristic Boundary): The gradient $\nabla_{\mathbf{U}}G_{i}$ is a left eigenvector of the Flux Jacobian $\mathbf{F}^{\prime}(\mathbf{U})$, or equivalently, the vector field points tangent to the boundary.
We verify these conditions for the defining constraints of $\Sigma$:
1. The Diffusion Condition: Since the artificial viscosity is scalar ($D=\varepsilon I_{2}$), every non-zero vector in $\mathbb{R}^{2}$ is a left eigenvector of $D$. Thus, this condition is trivially satisfied for the gradients of $w$, $z$, and $\rho$ everywhere on the boundary.
2. Convexity of $\Sigma$: The region $\Sigma$ is the intersection of three sets.
  • The constraint $\rho\geq\delta$ defines a convex half-plane in the $(\rho,m)$ coordinates.
[…]
    $m^{\prime\prime}(\rho)=-\frac{2\tilde{P}^{\prime}(\rho)+\rho\tilde{P}^{\prime\prime}(\rho)}{2\rho\tilde{c}(\rho)}.$
    Since our weighted pressure $\tilde{P}$ inherits the genuine nonlinearity condition $2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}>0$ for all $\rho>\delta$ (Lemma 2.1), we have $m^{\prime\prime}(\rho)<0$. This implies the boundary is concave and the sublevel set $\{w\leq C_{\sup}\}$ is convex. Similarly, for $z=\text{const}$, $m^{\prime\prime}(\rho)>0$, implying convexity of the superlevel set.
  • Thus, $\Sigma$ is the intersection of convex sets and is therefore convex.
3. The Flux Condition on Boundaries: We examine the invariance of the boundaries determined by the hyperbolic character of the system.
  • Upper and Left Boundaries ($w=C_{\sup}$ and $z=C_{\inf}$): By the definition of Riemann invariants, their gradients $\nabla_{\mathbf{U}}w$ and $\nabla_{\mathbf{U}}z$ are the left eigenvectors of the Flux Jacobian $\mathbf{F}^{\prime}(\mathbf{U})$ corresponding to eigenvalues $\lambda_{2}=u+\tilde{c}$ and $\lambda_{1}=u-\tilde{c}$, respectively. Thus, these boundaries are characteristic surfaces. The convective flux is tangent to the boundary $\partial\Sigma$, preventing the solution from crossing it.
  • The Lower Boundary ($\rho=\delta$): This boundary corresponds to the degeneracy locus $w(\mathbf{U})=z(\mathbf{U})$. We observe that as $\rho\to\delta$, the sound speed vanishes ($\tilde{c}\to 0$). Consequently, the two eigenvalues coalesce:
    $\lim_{\rho\to\delta}\lambda_{1}=\lim_{\rho\to\delta}\lambda_{2}=u.$
    Correspondingly, the gradients of the Riemann invariants become collinear. Since both $\nabla w$ and $\nabla z$ are left eigenvectors corresponding to the limit eigenvalue $u$, any linear combination of them is also a left eigenvector for $u$. Specifically, the gradient of the boundary defining function G=zwG=z-w satisfies:
    $\nabla_{\mathbf{U}}(z-w)\cdot\mathbf{F}^{\prime}(\mathbf{U})\Big|_{\rho=\delta}=(\nabla z\cdot\mathbf{F}^{\prime}-\nabla w\cdot\mathbf{F}^{\prime})=(u\nabla z-u\nabla w)=u\nabla_{\mathbf{U}}(z-w).$
    This proves that the boundary $\rho=\delta$ is a characteristic boundary with characteristic speed $u$. The flow is strictly tangential to this boundary in the phase space. Physically, this means the fluid velocity matches the boundary propagation speed, making it impossible for the particle trajectories to cross $\rho=\delta$.
Since all geometric conditions of the CCS theory are satisfied, $\Sigma$ is a positively invariant region. This establishes the uniform bounds independent of $\varepsilon$:
$\delta\leq\rho^{\varepsilon}(x,t)\leq M,\quad|u^{\varepsilon}(x,t)|\leq M.$
4.3 A Geometric Re-interpretation of the work of Lu
In this section, we provide a geometric verification of the invariant regions constructed in the seminal work of Lu [4]. While Lu originally established the existence of these regions via coordinate transformations and energy estimates, we demonstrate here that his result can be derived directly from the Chueh-Conley-Smoller (CCS) theory [8]. This analysis reveals that the modification of the mass flux in [4] serves precisely to satisfy the tangential flow condition required by the CCS framework.
4.3.1 Lu’s Modified System
Recall the regularization system proposed in [4]:
$\left\{\begin{aligned} &\rho_{t}+(\rho u-2\delta u)_{x}=\varepsilon\rho_{xx},\\ &(\rho u)_{t}+(\rho u^{2}-\delta u^{2}+P_{1}(\rho,\delta))_{x}=\varepsilon(\rho u)_{xx}.\end{aligned}\right$. (4.8)
The perturbed pressure is defined as $P_{1}(\rho,\delta)=\int_{2\delta}^{\rho}\frac{s-2\delta}{s}P^{\prime}(s)\,ds$. The invariant region proposed by Lu is:
$\Sigma^{Lu}=\left\{(\rho,u):\rho\geq 2\delta,\quad z^{Lu}\geq C_{1},\quad w^{Lu}\leq C_{2}\right\},$ (4.9)
where $z^{Lu},w^{Lu}$ are the Riemann invariants associated with the system (4.8).
4.3.2 Proof via CCS Theory
We now prove the invariance of $\Sigma^{Lu}$ using the geometric criteria of [8].
Theorem 4.3.
The region $\Sigma^{Lu}$ is positively invariant for the system (4.8) with respect to the CCS theory.
Proof.
Let $\mathbf{U}=(\rho,m)^{T}$. The system (4.8) takes the form $\mathbf{U}_{t}+\mathbf{F}_{Lu}(\mathbf{U})_{x}=\varepsilon\mathbf{U}_{xx}$. The diffusion matrix is scalar ($D=\varepsilon I$), satisfying the diffusion condition trivially. We focus on the Flux Condition at the boundaries of $\Sigma^{Lu}$.
1. The Hyperbolic Boundaries ($z^{Lu}=C_{1},w^{Lu}=C_{2}$): The system (4.8) is strictly hyperbolic for $\rho>2\delta$. By definition, the gradients of the Riemann invariants are the left eigenvectors of the Jacobian $\mathbf{F}_{Lu}^{\prime}(\mathbf{U})$. Thus, the convective vector field is tangential to the level sets of $z^{Lu}$ and $w^{Lu}$. The standard CCS condition for hyperbolic boundaries is satisfied.
2. The Artificial Boundary ($\rho=2\delta$): The core difficulty lies in the lower boundary $\rho=2\delta$. We must verify that the vector field $\mathbf{F}_{Lu}$ does not point outward across the surface $\rho=2\delta$. Let us examine the eigenvalues of Lu’s system. The modified flux Jacobian has eigenvalues (see Eq. (2.1) in [4]):
$\lambda_{1,2}^{Lu}=u\mp\frac{\rho-2\delta}{\rho}\sqrt{P^{\prime}(\rho)}.$ (4.10)
The crucial observation is the behavior at the boundary $\rho=2\delta$. Substituting $\rho=2\delta$ into (4.10):
$\lambda_{1}^{Lu}=\lambda_{2}^{Lu}=u.$ (4.11)
Geometrically, this implies that the characteristic curves become exactly tangent to the boundary $\rho=2\delta$ in the $(x,t)$ plane. In the phase space $(\rho,m)$, the boundary $\rho=2\delta$ corresponds to the degeneracy locus where the two families of characteristics coincide. Specifically, let $G(\mathbf{U})=2\delta-\rho\leq 0$. The normal vector is $\mathbf{n}=(-1,0)^{T}$. The condition that the flow is tangential means that the boundary propagates at the characteristic speed. Since $\lambda=u$, the fluid particles at the boundary move exactly with the boundary. To see this rigorously using the CCS condition, we verify that $\nabla\rho=(1,0)$ is a left eigenvector of $\mathbf{F}_{Lu}^{\prime}$ at $\rho=2\delta$ corresponding to the eigenvalue $u$. The mass conservation equation in (4.8) can be rewritten as:
$\rho_{t}+u\rho_{x}+(\rho-2\delta)u_{x}=\varepsilon\rho_{xx}.$
At the boundary where $\rho=2\delta$, the term $(\rho-2\delta)u_{x}$ vanishes. The equation degenerates to a pure convection-diffusion equation:
$\rho_{t}+u\rho_{x}=\varepsilon\rho_{xx}.$
This confirms that the convective transport velocity normal to the boundary $\rho=2\delta$ is exactly $u$. Since the boundary moves with the fluid, there is no cross-boundary flux. Thus, the condition $\nabla G\cdot\mathbf{F}_{Lu}^{\prime}\propto\nabla G$ holds, satisfying the CCS tangency requirement.
Therefore, the boundaries are either characteristic surfaces or degenerate characteristic surfaces where the flow is tangential, and the convexity of the region is guaranteed by the structure of gas dynamics, the region $\Sigma^{Lu}$ is positively invariant. ∎
Remark 4.4 (Geometric Unity with Lu’s Regularization).
It is illuminating to interpret the regularization strategy of Lu [4] and the current weighted pressure approach through the lens of the Chueh-Conley-Smoller theory. Although the two methods employ distinct mechanisms—Lu modifies the mass flux to $\rho u-2\delta u$, whereas we perturb the pressure to $\tilde{P}$—they share the identical geometric essence.
From the perspective of invariant regions, the core requirement is that the boundary of the domain must be a characteristic surface where the vector field is tangential (the Flux Condition).
  • Lu’s Approach: The flux modification in [4] is algebraically designed to force the eigenvalues to degenerate to $\lambda=u$ at the boundary $\rho=2\delta$, thereby satisfying the tangential condition via transport retardation.
  • Our Approach: We achieve the same characteristic degeneracy $\lambda=u\pm\tilde{c}\to u$ at $\rho=\delta$ by geometrically flattening the pressure such that the sound speed vanishes ($\tilde{c}\to 0$).
Thus, both methods represent different pathways to satisfy the same geometric constraint required for the existence of invariant regions. However, our approach distinguishes itself by satisfying this geometric condition while strictly preserving the physical law of mass conservation.
@@ 4.4 Energy Dissipation and $H^{-1}$ Compactness
$D^{\varepsilon}:=\varepsilon(\mathbf{U}^{\varepsilon}_{x})^{T}\nabla^{2}\eta^{*}\mathbf{U}^{\varepsilon}_{x}=\varepsilon\Big(\frac{\tilde{P}^{\prime}(\rho^{\varepsilon})}{\rho^{\varepsilon}}(\rho^{\varepsilon}_{x})^{2}+\rho^{\varepsilon}(u^{\varepsilon}_{x})^{2}\Big)\geq 0.$